Seventy-two families were asked how many pets they have with the following results. Find the Mean, Median, and Mode of the data set. Show your work for the Mean. MEAN: (Show your work.) MEDIAN: (Show your work.) MODE:

Number of pets Number of families
0 10
1 20
2 22
3 12
4 6
5 2
Total 72
Answer #1

Solution-a

Number of pets(x) Number of families(f) fx
0 10 0
1 20 20
2 22 44
3 12 36
4 6 24
5 2 10
total 134

Mean=\sum fx/\sum f

=134/72

=1.8611

Solution-b:

Number of pets(x) Number of families(f) Cumulative frequency
0 10 10
1 20 30
2 22 52 n/2=72/2=36
3 12 64
4 6 70
5 2 72
Total 72

median is n/2 th item

=36 th item

From cumulative frequency ,36 th item is 2

median=2

Solution-c

:

mode is highest frequency

Here highest frequency is 22

so mode is 2

mode =2